# Unconventional tunnel magnetoresistance scaling with altermagnets

### v1

@@ Abstract
In conventional magnetic tunnel junctions (MTJs), the tunnel magnetoresistance (TMR) typically increases with barrier thickness as electron transmission in the antiparallel configuration decays faster than that of the parallel configuration. In this work, we reveal an anomalous scaling effect in altermagnetic tunnel junctions (AMTJs), where the TMR decreases anomalously with an increasing barrier thickness. The anomalous scaling originates from the overlapping spin-split branches form a transmission path that cannot be suppressed in the antiparallel state. Such phenomena is explained by a double-barrier model and is further demonstrated using ab initio quantum transport simulations in 2D V2Te2O/Cr2Se2O/V2Te2O-based AMTJ, where the TMR anomalously decreases from 220% to 40% as the layer number of Cr2Se2O increases from 1 to 5. Our work identifies a peculiar unexpected transport characteristic of AMTJ, providing a fundamental limit on AMTJ device design and illustrating the potential optimal design of AMTJ at the ultrascaled monolayer limit.
@@
Introduction. Magnetic tunnel junctions (MTJs) are a fundamental component of spintronics, which constitute a nonvolatile memory function through the giant magnetoresistance effect [1, 2, 3, 4, 5, 6, 7, 8, 9]. The scaling of MTJs is a key trend in advancement of spintronics to achieve high-density storage, high-speed operation, and efficient spin orbital torque (SOT) [10, 11, 12, 13, 14, 15, 16]. Ferromagnetic memory faces difficulties to achieve higher storage density and device miniaturization due to the presence of stray fields [17, 18, 19, 20]. Antiferromagnetic tunnel junctions have gained attention for their robustness against magnetic perturbations, absence of stray fields, high Néel temperature, and ultrafast spin dynamics [21, 22, 23, 24, 25, 26]. However, conventional antiferromagnets exhibit degenerate energy bands, with only hidden spin polarization available for control. Altermagnets (AMs), featuring alternating spin polarizations in both real space and momentum space, have recently been recognized as a type of unconventional magnets [27, 28, 29, 30, 31, 32]. They exhibit a giant spin-splitting without a net magnetic moment, thus offering an ideal platform for ultracompact integration of MTJs [33].
@@
where and denote the conductance of the device in the parallel (P) and anti-parallel (AP) magnetic configurations, respectively [34]. In ferromagnetic and A-type antiferromagnetic tunnel junctions, TMR typically increases as the barrier layer thickness increases [35, 36, 37, 38, 39, 40, 41]. For example, in graphite/2D CrPS4/graphite MTJs, the TMR ratio rises by approximately when the number of CrPS4 layers is increased from 2 to 7 [42]. A similar trend is observed in Ag/CrI3/Ag MTJs, where the TMR ratio grows from 103% to 109% with the barrier thickness of CrI3 varying from 2 to 12 layers [37]. This behavior can be explained by the different decay rates of conductance in the P and AP configurations as described by the Julliere model [38, 43, 44, 34]. As shown in Fig. 1(a), the conductance is dominated by Bloch states with a small decay rate in P state, which is also called the majority spin conductance. In the AP configuration, the conductance is strongly suppressed due to spin mismatch between the electrodes, and it is mainly contributed by interface resonance states with large decay rate. As a result, TMR decreases as the barrier becomes thinner according to , where represents the difference in the decay rate between the AP and the P configurations.
@@
Anomalous scaling in AMTJs. In this work, we reveal an anomalous scaling behavior in altermagnetic tunnel junctions (AMTJs) [Fig. 1(b)], whose TMR decreases as the barrier thickness increases. This anomalous scaling originates from the distinct morphology of the altermagnetic Fermi surface, where the momentum-space overlap between the spin-split branches forms a persistent transmission channel that cannot be suppressed in the AP configuration. The states that suppress the increase in TMR are, in fact, the spin-degenerate Bloch states in altermagnets. Such spin degeneracy is general in altermagnets and protected by the symmetry when the little group of momentum includes at least one real-space symmetry operation that connects the opposite-spin sublattices. [29, 33]. In contrast, ferromagnetic or A-type antiferromagnetic MTJs as discussed above possess fully spin-split bands without overlap, allowing efficient suppression of transmission in the AP configuration. To illustrate the anomalous scaling, we develop a double-barrier model, which sub-divides the spin-split Fermi surfaces into an outer region and a central region, labelled as A and B in Figs. 1(c) and 1(d), respectively. In the AP configuration, Neel vectors of the incident and transmission leads are misaligned. Electron transmission mediated by states in Region A can be effectively switched off due to spin-mismatch in this case. However, transport channel in Region B remains open even in the AP configuration. This arises because of the overlapping of two spin-split Fermi surfaces which create a spin-degenerate transmission highway that remains open in both P and AP states [Figs. 1(c) and 1(d)]. As the barrier thickness increases, electron tunneling through the spin-degenerate highway (provided by the electronic states in Region B) remains high with higher OFF state current. Consequently, TMR reduces with increasing barrier thickness, leading to the anomalous scaling behavior as compared to the opposite behavior in conventional MTJs.
@@
Anomalous scaling mechanism. According to the tunneling theory, electron transmission decays exponentially as where is the barrier thickness [45]. Here is the imaginary part of the wavevector (), which can be written as
@@
where is the wavevector parallel to the tunneling interface, is the Fermi level [38, 43], and potential barrier is a function of and spin . We denote two different tunneling barriers for spin-degenerate transport mediated by states in Region B and spin-nondegenerate transport mediated by states in Region A as and , respectively. In AMTJ, the transport channel associated with is always open regardless of the P and AP configurations. In contrast, the transport channel through can be more effectively switched ON or OFF, thus playing the key role in determining the scaling of TMR. The total conductance under is composed of
@@
where is quantum conductance ( and are elementary charge and Planck constant, respectively) and , and the index corresponds to the contribution of electron transmission mediated through Region A and B, respectively. The total conductance in the P configuration is . In the AP configuration, the tunneling through interface resonance states in Region B can be neglected, resulting in a conductance of . The -related TMR ratio is
@@
The total TMR is the summation . Here we consider two scenarios: (i) If , the TMR increases with increasing barrier thickness , which is consistent with the conventional scaling behavior. (ii) If , which generally occurs in AMTJ, the conductance decays more rapidly than as increases as shown in Fig. 1(b). As a result, the TMR increases with decreasing , as shown at the bottom of Fig. 1(b), which corresponds to the anomalous scaling effect. We next demonstrate this anomalous scaling behavior in a representative AMTJ, corresponding to the second case discussed above.
@@
First principle calculation. We calculate the scaling behavior of 2D AMTJ based on 2D AM metal V2Te2O (VTO) and AM semiconductor Cr2Se2O (CSO). VTO and CSO are both room-temperature AM with high Néel temperature of 740 K and 350 K, respectively. The lattice mismatch of VTO/CSO heterojunction is less than 0.6% [46, 47]. The atomic structure of these two materials exhibits , , , and symmetry [Fig. 2(a)]. The spin polarization of Cr or V atoms are opposite in the unit cell [bottom right in Fig. 2(a)]. With such magnetic ordering, symmetry (where represents a half-lattice translation along [110]) is broken, leading to unconventional spin splitting in reciprocal space. This phenomenon is further validated by the density functional theory (DFT)-calculated monolayer (ML) band structures [Figs. 2(b) and (d)]. The -resolved spin-splitting energy at the conduction band minimum (CBM) of CSO and VTO [Figs. 2(c) and (e)] indicate that they both possess -wave anisotropy with spin degeneration around point. This band structure features a spin-degenerate region near the point and spin-split regions away from it, corresponding to Region B and Region A in the aforementioned mechanism. It is worth noting that this condition is not limited to the present system; similar characteristics can also be found in altermagnets with -wave or -wave anisotropy. The CSO is serving as an insulating barrier with a bandgap of 0.97 eV. The CSO in the insulating layer is AB-stacked, and its structure is shown in Fig. S2 in Supplementary Material [48]. To ensure consistent spin splitting in -space across different layers, an interlayer out-of-plane antiferromagnetic (AFM) spin alignment is adopted within the same material, forming a G-type configuration [49]. We construct a tight-binding (TB) model to describe the spin splitting characteristics of the CSO, which is consistent with the results from DFT calculations (Fig. S1 in Supplementary Material [48]).
@@
We perform DFT calculations coupled with the nonequilibrium Green’s function (NEGF) method, as implemented in QuantumATK (see Supplementary Materials for computational details [48]) [52], to assess the transport properties of the AMTJ. In the VTO/1L-CSO/VTO AMTJ, switching from the AP state yields TMR ratios of about 220% for the P state and 137% for the I state. The distinct transport behaviors between P and I configurations can be attributed to the spin-dependent tunneling barrier. The difference in transmission between the P and I configurations arises from the lower tunneling barrier in the P state, where the spin-splitting of the electrodes and barrier are mutually aligned (Figs. 4e,f). Likewise, the transmission eigenstates for different spin channels further reveal that the tunneling barrier due to the spin-splitting mismatch between CSO and VTO effectively suppresses electron transmission in the I state (see Figs. S4 in Supplementary Material [48]). The lowest transmission observed in the AP state (Fig. 4g) indicates that spin matching between electrodes plays a more dominant role in tunneling than spin alignment within the barrier. Especially, spin-degenerate states (near the point) consistently exhibit lower tunneling barriers due to the constant spin alignment in all three magnetic configurations and dominate the transport in the AP state.
@@
Next, we investigate the tunneling of AMTJ with varying CSO layer number. With increasing CSO thickness, all three magnetic configurations exhibit exponential decay in transmission. However, the decay rates differ significantly: the AP configuration shows the slowest decay (only 31%), in contrast to 72.7% and 59.3% reductions in the P and I states, respectively. This trend highlights a persistent transmission channel in the AP configuration that is less sensitive to barrier thickness. As a result, the TMR drops anomalously—from 220% to 40% and from 137% to 39% for P-to-AP and I-to-AP switching, respectively—as the number of CSO layers increases from 1 to 5 (Fig. 5a). The detailed energy-dependent transmissions for different magnetic configurations are shown in Fig. S3(b)–(c) (Supplementary Material [48]).
@@
In summary, we reveal an anomalous scaling behavior of decreasing TMR with increasing barrier thickness in AMTJ, which deviates from conventional expectations. The underlying mechanism of this anomalous scaling lies in the presence of a persistent transmission channel arising from spin-degenerate states, which remains largely unaffected by switching between the P and AP configurations. To qualitatively interpret this behavior, we further construct a simplified double-barrier model that distinguishes between effective and ineffective spin-selective transmission. The effective part, originating from spin-split states, decays rapidly with increasing barrier thickness. While the ineffective part exhibits weak decay and ultimately dominates the overall transmission at larger thicknesses. This contrasts with conventional ferromagnetic MTJs, where the bandstructure is fully split and only the effective spin-selective transmission is present. To validate this mechanism, we conduct ab initio quantum transport simulations on a V2Te2O/Cr2Se2O/V2Te2O AMTJ, demonstrating that increasing the Cr2Se2O layer thickness from 1 to 5 layers led to a significant TMR degradation from 220% to 40%. Our results not only provide new insights into the fundamental transport behavior of altermagnets but also highlight a key design consideration fundamentally impacting the performance of magnetic tunneling junction based on altermagnets.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 12274002 and 91964101), the Ministry of Science and Technology of China (No. 2022YFA1203904), the Fundamental Research Funds for the Central Universities, the High-performance Computing Platform of Peking University and the MatCloud + high-throughput materials simulation engine, and the National Research Foundation, Singapore, under its Frontier Competitive Research Programme (NRF-F-CRP-2024-0001). L. K. A. is supported by the Singapore ASTAR IRG (M23M6c0102).References
@@
- [48] S. Material, Supplementary material, .
@@
- Zhang et al. [2025] F. Zhang, X. Cheng, Z. Yin, C. Liu, L. Deng, Y. Qiao, Z. Shi, S. Zhang, J. Lin, Z. Liu, M. Ye, Y. Huang, X. Meng, C. Zhang, T. Okuda, K. Shimada, S. Cui, Y. Zhao, G.-H. Cao, S. Qiao, J. Liu, and C. Chen, Crystal-symmetry-paired spin–valley locking in a layered room-temperature metallic altermagnet candidate, Nat. Phys. (2025).

### v2

@@ Abstract
In conventional magnetic tunnel junctions (MTJs), the tunnel magnetoresistance (TMR) typically increases with barrier thickness as electron transmission in the antiparallel configuration decays faster than that of the parallel configuration. In this work, we reveal an anomalous scaling effect in altermagnetic tunnel junctions (AMTJs), where the TMR decreases anomalously with an increasing barrier thickness. The anomalous scaling originates from the overlapping spin-split branches forming a transmission path that cannot be suppressed in the antiparallel state. Such phenomenon is explained by a double-barrier model and is further demonstrated using ab initio quantum transport simulations in 2D V2Te2O/Cr2Se2O/V2Te2O-based AMTJ, where the TMR anomalously decreases from 220% to 40% as the layer number of Cr2Se2O increases from 1 to 5. Our work identifies a peculiar unexpected transport characteristic of AMTJ, providing a fundamental limit on AMTJ device design and illustrating the potential optimal design of AMTJ at the ultrascaled monolayer limit.
@@
Introduction. Magnetic tunnel junctions (MTJs) are a fundamental component of spintronics, which constitute a nonvolatile memory function through the giant magnetoresistance effect [1, 2, 3, 4, 5, 6, 7, 8, 9]. The scaling of MTJs is a key trend in the advancement of spintronics to achieve high-density storage, high-speed operation, and efficient spin–orbit torque (SOT) [10, 11, 12, 13, 14, 15, 16]. Ferromagnetic memory faces difficulties to achieve higher storage density and device miniaturization due to the presence of stray fields [17, 18, 19, 20]. Antiferromagnetic tunnel junctions have gained attention for their robustness against magnetic perturbations, absence of stray fields, high Néel temperature, and ultrafast spin dynamics [21, 22, 23, 24, 25, 26]. However, conventional antiferromagnets exhibit degenerate energy bands, with only hidden spin polarization available for control. Altermagnets (AMs), featuring alternating spin polarizations in both real space and momentum space, have recently been recognized as a type of unconventional magnets [27, 28, 29, 30, 31, 32, 33, 34, 35]. They exhibit a giant spin-splitting without a net magnetic moment, thus offering an ideal platform for ultracompact integration of MTJs [36].
@@
where and denote the conductance of the device in the parallel (P) and antiparallel (AP) magnetic configurations, respectively [37]. In ferromagnetic and A-type antiferromagnetic tunnel junctions, TMR typically increases as the barrier layer thickness increases [38, 39, 40, 41, 42, 43, 44]. For example, in graphite/2D CrPS4/graphite MTJs, the TMR ratio rises by approximately when the number of CrPS4 layers is increased from 2 to 7 [45]. A similar trend is observed in Ag/CrI3/Ag MTJs, where the TMR ratio grows from 103% to 109% with the barrier thickness of CrI3 varying from 2 to 12 layers [40]. This behavior can be explained by the different decay rates of conductance in the P and AP configurations as described by the Julliere model [41, 46, 47, 37]. As shown in Fig. 1(a), the conductance is dominated by Bloch states with a small decay rate in P state, which is also called the majority spin conductance. In the AP configuration, the conductance is strongly suppressed due to spin mismatch between the electrodes, and it is mainly contributed by interface resonance states with a large decay rate. As a result, TMR decreases as the barrier becomes thinner according to , where represents the difference in the decay rate between the AP and the P configurations.
@@
Anomalous scaling in AMTJs. In this work, we reveal an anomalous scaling behavior in altermagnetic tunnel junctions (AMTJs) [Fig. 1(b)], whose TMR decreases as the barrier thickness increases. This anomalous scaling originates from the distinct morphology of the altermagnetic Fermi surface, where the momentum-space overlap between the spin-split branches forms a persistent transmission channel that cannot be suppressed in the AP configuration. The states that suppress the increase in TMR are, in fact, the spin-degenerate Bloch states in altermagnets. Such spin degeneracy is general in altermagnets and protected by the symmetry when the little group of momentum includes at least one real-space symmetry operation that connects the opposite-spin sublattices. [29, 36]. In contrast, ferromagnetic or A-type antiferromagnetic MTJs as discussed above possess fully spin-split bands without overlap, allowing efficient suppression of transmission in the AP configuration. To illustrate the anomalous scaling, we develop a double-barrier model, which subdivides the spin-split Fermi surfaces into an outer region and a central region, labeled A and B in Figs. 1(c) and 1(d), respectively. In the AP configuration, Néel vectors of the incident and transmission leads are misaligned. Electron transmission mediated by states in Region A can be effectively switched off due to spin mismatch in this case. However, the transport channel in Region B remains open even in the AP configuration. This arises because of the overlap of two spin-split Fermi surfaces which create a spin-degenerate transmission highway that remains open in both P and AP states [Figs. 1(c) and 1(d)]. As the barrier thickness increases, electron tunneling through the spin-degenerate highway (provided by the electronic states in Region B) remains high with higher OFF state current. Consequently, TMR reduces with increasing barrier thickness, leading to the anomalous scaling behavior as compared to the opposite behavior in conventional MTJs.
@@
Anomalous scaling mechanism. According to tunneling theory, the transmission of electrons decays exponentially as where is the barrier thickness [48] . Here is the imaginary part of the wavevector (), which can be written as
@@
where is the wavevector parallel to the tunneling interface, is the Fermi level [41, 46], and potential barrier is a function of and spin . Here, the density of states of the initial and final states are uniform in -space. We denote two different tunneling barriers for spin-degenerate transport mediated by states in Region B and spin-nondegenerate transport mediated by states in Region A as and , respectively. In AMTJ, the transport channel associated with is always open regardless of the P and AP configurations. In contrast, the transport channel through can be more effectively switched ON or OFF, thus playing the key role in determining the scaling of TMR. The total conductance under is composed of
@@
where is quantum conductance ( and are elementary charge and Planck constant, respectively) and , and the index corresponds to the contribution of electron transmission mediated through Region A and B, respectively. The total conductance in the P configuration is . In the AP configuration, the tunneling through interface resonance states in Region B can be neglected, resulting in a conductance of . The TMR ratio is
| TMR | ||||
@@
Here we consider two scenarios: (i) If , the TMR increases with increasing barrier thickness , which is consistent with the conventional scaling behavior. (ii) If , the conductance decays more rapidly than as increases as shown in Fig. 1(b). As a result, the TMR increases with decreasing , as shown at the bottom of Fig. 1(b), which corresponds to the anomalous scaling effect. While the relative -space area of regions A and B does not affect the monotonicity of the TMR trend, it will influence the overall decay rate. We next demonstrate this anomalous scaling behavior in a representative AMTJ, corresponding to the second case discussed above.
@@
First principle calculation. We calculate the scaling behavior of 2D AMTJ based on 2D AM metal V2Te2O (VTO) and AM semiconductor Cr2Se2O (CSO). VTO and CSO are both room temperature AM with high Néel temperature of 740 K and 350 K, respectively. The lattice mismatch of VTO/CSO heterojunction is less than 0.6% [49, 50]. The presence of spin-orbital coupling (SOC) can affect both the band structure and spin hybridization of a system. Therefore, we also investigated the impact of SOC on VTO and CSO (see Supplementary Fig. S1 [51]). We found that SOC has a negligible effect on their band structures which is also illustrated in previous studies [49, 50]. Moreover, due to the protection from spin-mirror coupling, the spin hybridization in VTO and CSO is significantly suppressed [52]. Therefore, SOC is not included in the subsequent calculations. The atomic structure of these two materials exhibits , , , and symmetry [Fig. 2(a)]. The spin polarization of neighboring Cr or V atoms are opposite in the unit cell [bottom right in Fig. 2(a)]. With such magnetic ordering, symmetry (where represents a half-lattice translation along [110]) is broken, leading to unconventional spin splitting in reciprocal space. This phenomenon is further validated by the density functional theory (DFT)-calculated monolayer (ML) band structures [Figs. 2(b) and (d)]. The -resolved spin-splitting energy at the conduction band minimum (CBM) of CSO and VTO [Figs. 2(c) and (e)] indicate that they both possess -wave anisotropy with spin degeneration around point. This band structure features a spin-degenerate region near the point and spin-split regions away from it, corresponding to Region B and Region A in the aforementioned mechanism. The CSO is serving as an insulating barrier with a bandgap of 0.97 eV. The CSO in the insulating layer is AB-stacked, and its structure is shown in Fig. S3 in Supplementary Material [51]. To ensure consistent spin splitting in -space across different layers, an interlayer out-of-plane antiferromagnetic (AFM) spin alignment is adopted within the same material, forming a G-type configuration [53]. We construct a tight-binding (TB) model to describe the spin splitting characteristics of the CSO, which is consistent with the results from DFT calculations (Fig. S2 in Supplementary Material [51]). Nearly identical total energies of both the individual CSO and VTO layers, as well as the CSO/VTO heterostructure under antiferromagnetic (G-type) and ferromagnetic (C-type) configuration, suggest that the original altermagnetic order of each layer is preserved after stacking (Table S1 in Supplementary Material [51]).
@@
We perform DFT calculations coupled with the nonequilibrium Green’s function (NEGF) method, as implemented in QuantumATK (see Supplementary Materials for computational details [51]) [56], to assess the transport properties of the AMTJ. In the VTO/1L-CSO/VTO AMTJ, switching from the AP state yields TMR ratios of about 220% for the P state and 137% for the I state. The difference in transport behaviors between P and I configurations can be attributed to the spin-dependent tunneling barrier. The difference in transmission between the P and I configurations arises from the lower tunneling barrier in the P state, where the spin-splitting of the electrodes and barrier are mutually aligned (Figs. 4e,f). Likewise, the transmission eigenstates for different spin channels further reveal that the tunneling barrier due to the spin-splitting mismatch between CSO and VTO effectively suppresses electron transmission in the I state (see Figs. S4 in Supplementary Material [51]). The lowest transmission observed in the AP state (Fig. 4g) indicates that spin matching between electrodes plays a more dominant role in tunneling than spin alignment within the barrier. Especially, spin-degenerate states (near the point) consistently exhibit lower tunneling barriers due to the constant spin alignment in all three magnetic configurations and dominate the transport in the AP state.
@@
Next, we investigate the tunneling of AMTJ with varying CSO layer number. With increasing CSO thickness, all three magnetic configurations exhibit exponential decay in transmission. However, the decay rates differ significantly: the AP configuration shows the slowest decay (only 31%), in contrast to 72.7% and 59.3% reductions in the P and I states, respectively. This trend highlights a persistent transmission channel in the AP configuration that is less sensitive to barrier thickness. As a result, the TMR drops anomalously: from 220% to 40% and from 137% to 39% for the P-to-AP and I-to-AP switching, respectively, as the number of CSO layers increases from 1 to 5 (Fig. 5a). Theoretically, TMR continues to decay with increasing barrier thickness and asymptotically approaches zero. In practice, however, as the tunneling current diminishes with thickness, it may eventually fall below the detection limit of experimental setups. These considerations imply two distinct critical thicknesses: one defined by the experimental measurability of the current, and the other by the theoretical vanishing of TMR. The relative importance of each depends on the specific material parameters and measurement conditions. The detailed energy-dependent transmissions for different magnetic configurations are shown in Fig. S4(b)–(c) (Supplementary Material [51]).
Discussion and conclusion. Although our study focuses on altermagnet-based tunnel junctions, the applicability of our theoretical framework is may potentially be extended to a wide range of momentum–spin coupled systems, such as spin–orbit coupling (SOC) [57, 58], non-collinear antiferromagnets [59], and spin–valley locking [60], etc. The key point of this effect lies in the fact that spin-degenerate and non-degenerate Bloch waves exhibit different decay rates. In addition, pinholes or resonant tunneling in conventional antiferromagnetic devices may also lead to anomalous scaling behavior [61]. This anomalous scaling behavior in TMR does not necessarily occur only in altermagnets and therefore cannot be used as definitive evidence for identifying altermagnetic materials.
In addition, to further examine the robustness of this effect and address possible experimental realizations, we have also considered an alternative junction geometry with altermagnetic electrodes and a nonmagnetic barrier. Using ZnSe as a representative 2D semiconductor barrier, we find a similar decay of TMR with increasing barrier thickness (See Fig. S6 of the Supplementary Material for details [51]). This confirms that the unconventional TMR scaling is not restricted to the all-altermagnetic configuration but can also emerge in more experimentally accessible structures.
@@
In summary, we reveal an anomalous scaling behavior of decreasing TMR with increasing barrier thickness in AMTJ, which deviates from conventional expectations. The underlying mechanism of this anomalous scaling lies in the presence of a persistent transmission channel arising from spin-degenerate states, which remains largely unaffected by switching between the P and AP configurations. To qualitatively interpret this behavior, we further construct a simplified double-barrier model that distinguishes between effective and ineffective spin-selective transmission. The effective part, originating from spin-split states, decays rapidly with increasing barrier thickness. While the ineffective part exhibits weak decay and ultimately dominates the overall transmission at larger thicknesses. Whether the TMR exhibits an unconventional scaling behavior is primarily determined by the difference in decay constants between the ineffective and effective regions. The overall decay rate of TMR with barrier thickness is further influenced by both the relative area ratio of these -space regions and their respective decay rates. We conduct ab initio quantum transport simulations on a V2Te2O/Cr2Se2O/V2Te2O AMTJ, demonstrating that increasing the Cr2Se2O layer thickness from 1 to 5 layers led to a significant TMR degradation from 220% to 40%. Our results not only provide new insights into the fundamental transport behavior of altermagnets but also highlight a key design consideration that fundamentally impacts the performance of the magnetic tunnel junction based on altermagnets.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 12274002 and 91964101), the Ministry of Science and Technology of China (No. 2022YFA1203904), the China Scholarship Council (CSC), the Fundamental Research Funds for the Central Universities, the High-performance Computing Platform of Peking University and the National Research Foundation, Singapore, under its Frontier Competitive Research Programme (NRF-F-CRP-2024-0001). L. K. A. is supported by the Singapore ASTAR IRG (M23M6c0102). We thank Dingfu Shao for insightful discussions during 2025 Workshop on Altermagnets and School on Magnetic Symmetry (Shanghai). The authors also thank Dr. Yuntian Liu for helpful discussions.References
@@
- Fu et al. [2025a] P.-H. Fu, S. Mondal, J.-F. Liu, Y. Tanaka, and J. Cayao, Floquet engineering spin triplet states in unconventional magnets (2025a), arXiv:2505.20205 .
- Fu et al. [2025b] P.-H. Fu, Q. Lv, Y. Xu, J. Cayao, J.-F. Liu, and X.-L. Yu, All-electrically controlled spintronics in altermagnetic heterostructures (2025b), arXiv:2506.05504 .
- Fukaya et al. [2025] Y. Fukaya, B. Lu, K. Yada, Y. Tanaka, and J. Cayao, Superconducting phenomena in systems with unconventional magnets (2025), arXiv:2502.15400 .
@@
- [51] Supplementary material, .
- Zhang et al. [2025a] R.-W. Zhang, C. Cui, Y. Wang, J. Duan, Z.-M. Yu, and Y. Yao, Quantized spin-hall conductivity in altermagnet Fe2Te2O with mirror-spin coupling (2025a), arXiv:2503.10681 .
@@
- Huang et al. [2025] K. Huang, K. Samanta, D.-F. Shao, and E. Y. Tsymbal, Two-dimensional nonvolatile valley spin valve, ACS Nano 19, 3448 (2025).
- Tao et al. [2025] L. Tao, M. Dou, X. Wang, and E. Tsymbal, Ferroelectric spin-orbit valve effect, Phys. Rev. Lett. 134, 076801 (2025).
- Rimmler et al. [2025] B. H. Rimmler, B. Pal, and S. S. P. Parkin, Non-collinear antiferromagnetic spintronics, Nat Rev Mater 10, 109 (2025).
- Yan et al. [2025] K. Yan, L. Cheng, Y. Hu, J. Gao, X. Zou, and X. Chen, Giant tunneling magnetoresistance based on spin-valley-mismatched ferromagnetic metals, Phys. Rev. Lett. 134, 036302 (2025).
- Suzuki and Mizukami [2023] K. Z. Suzuki and S. Mizukami, Tunnel magnetoresistance exceeding 100% in magnetic tunnel junctions using mn-based tetragonal alloy electrodes with perpendicular magnetic anisotropy, AIP Advances 13, 035225 (2023).
@@
- Zhang et al. [2025b] F. Zhang, X. Cheng, Z. Yin, C. Liu, L. Deng, Y. Qiao, Z. Shi, S. Zhang, J. Lin, Z. Liu, M. Ye, Y. Huang, X. Meng, C. Zhang, T. Okuda, K. Shimada, S. Cui, Y. Zhao, G.-H. Cao, S. Qiao, J. Liu, and C. Chen, Crystal-symmetry-paired spin–valley locking in a layered room-temperature metallic altermagnet candidate, Nat. Phys. (2025b).